Taking 20 ml of this solution, 50 ml of pure water is added on it and it is titrated with 4 M NaOH solution. At the turning point, what is the concentration of the H2SO4 solution used, since 35 ml of the NaOH solution is spent?
1
Expert's answer
2020-12-28T04:41:39-0500

no. of milli eq. of NaOH used = 4× 35 = 140 milli eq.

so that much milli eq. of H2SO4 used.


So concentration of H2SO4 = 140/70 = 2 Normal


Molarity = 2/(n-factor) = 2/2 Molar= 1 Molar